Demand, Supply and Equilibrium: Given the following equations representing the behavior of producers and consumers:...

  1. Demand, Supply and Equilibrium:

Given the following equations representing the behavior of producers and consumers:

Price

Quantity Demanded Qd

Quantity Supplied Qs

52

48

44

40

35

32

29

26

                    24                                                                                                        

Consumers: Qd = 3,380 - 35P, Producers: Qs =95P, (P: Price) (Qd: quantity demanded, Qs: Quantity supplied )

  1. What price corresponds to the equilibrium price for this market? (1%) What is the equilibrium quantity?   
  2. Over what range of prices does a Surplus result?
  3. Over what range of prices does a Shortage result?
  4. If a surplus exists, explain the process by which market prices will adjust? (2%) Will these prices adjust upwards or downwards?
  5. Please graph the above price and quantity data and carefully label the Supply curve, demand curve, Equilibrium price, and Equilibrium quantity.
Answer #1

As per rules and time constraint, only one question with max four sub parts is allowed. Kindly ask last part separately.

Ans) At equilibrium, quantity demanded is equal to quantity supplied. This is market clearing point.

When price is above equilibrium price, there is surplus or excess supply in the market. Here, there is downward pressure on the price. Price will decrease till it reaches equilibrium.

When price is below equilibrium price, there is shortage or excess demand. Here,there is upward pressure on the prices. Price will increase till it reaches equilibrium.

Price above equilibrium price Equilibrium Price below equilibrium / price

Q = 95P Qs = 95C 52) Qs = 4,940 Os- 95 (48) Q = 4,560 5- CSURPLUS) AT P: 52 $1 = 3,380-35P Q = 3,380 – 35 (52) Py = 1560 (SUR

AT P= 32 0 = 3380- ©5732 Qd=2260 AT P=ga Qd = 3080 - 35 (29) Go = 2,565 Os= 95 (32) Qs = 3040 CSURPLUS) Qs = 95 (29) Qs = 2,7